Find the Wronskian of the following functions and determine whether it is linearly dependent or linearly independent on (-∞,∞).

  1. {x2, x+1, x-3}                ans, W=8, linearly independent
  2. {3e2x, e2x}                     ans, W=0, linearly dependent
  3. {x2, x3, x4}                    ans, W=2x^6, linearly independent
1
Expert's answer
2021-12-23T18:08:15-0500

1.



"W(y_1, y_2, y_3, x)=\begin{vmatrix}
 x^2 & x+1 & x-3 \\
 2x & 1 & 1 \\
 2 & 0 & 0
\end{vmatrix}""=2\begin{vmatrix}
 x+1 & x-3 \\
 1 & 1
\end{vmatrix}=2(x+1-x+3)=8\not=0"

Therefore, the set "\{x^2, x+1, x-3\}" is linearly independent on "(-\infin, \infin)."


2.



"W(y_1, y_2, x)=\begin{vmatrix}
 3e^{2x} & e^{2x} \\
 6e^{2x} & 2e^{2x}
\end{vmatrix}""=6e^{2x}-6e^{2x}=0"

Therefore, the set "\{3e^{2x}, e^{2x}\}" is linearly dependent on "(-\infin, \infin)."


3.



"W(y_1, y_2, y_3, x)=\begin{vmatrix}
 x^2 & x^3 & x^4 \\
 2x & 3x^2 & 4x^3 \\
 2 & 6x & 12x^2
\end{vmatrix}""=x^2(36x^4-24x^4)-x^3(24x^3-8x^3)""+x^4(12x^2-6x^2)=12x^6-16x^6+6x^6""=2x^6\not=0, except\ at\ \ x=0"

Therefore, the set "\{x^2, x^3, x^4\}" is linearly independent on "(-\infin, \infin)."